the length of the garden is six feet more than three times the width. eighty four feet of total fencing is required .whats the length and width of the garden?
Answer by checkley77(12844)

L=3W+6
2L+2W=84
2(3W+6)+2W=84
6W+12+2W=84
8W=84-12
8W=72
W=72/8
W=9 FT. ANS. FOR THE WIDTH.
L=3*9+6
L=27+6
L=33 FT. ANS. FOR THE LENGTH.
PROOF:
2*33+2*9=84
66+18=84
84=84